OpenStudy (anonymous):

Find the slope of the line passing through (9, -6) and (5, 10)

OpenStudy (anonymous):

There is no way your just curious! slope: y2 - y1 / x2 - x1

OpenStudy (anonymous):

yeah there is stop answering my questions @jazy

OpenStudy (anonymous):

and how would you write that

OpenStudy (anonymous):

answer please

OpenStudy (anonymous):

10 -(-6) = 16 5 - 9 = -4 16/-4 simplify and you have your slope

OpenStudy (anonymous):

-4 is what I got

OpenStudy (anonymous):

^^right:)

OpenStudy (anonymous):

-4?

OpenStudy (anonymous):

yes: 16/-4 = -4/1 = -4

OpenStudy (anonymous):